# Networks with many structural scales: a Renormalization Group perspective

### v2

@@ Abstract
Scale invariance profoundly influences the dynamics and structure of complex systems, spanning from critical phenomena to network architecture. Here, we propose a precise definition of scale-invariant networks by leveraging the concept of a constant entropy loss rate across scales in a renormalization-group coarse-graining setting. This framework enables us to differentiate between scale-free and scale-invariant networks, revealing distinct characteristics within each class. Furthermore, we offer a comprehensive inventory of genuinely scale-invariant networks, both natural and artificially constructed, demonstrating, e.g., that the human connectome exhibits notable features of scale invariance. Our findings open new avenues for exploring the scale-invariant structural properties crucial in biological and socio-technological systems.
@@
The network paradigm effectively captures essential attributes of real-world complex systems, offering a natural framework for studying entangled interconnected systems across disciplines like neuroscience Fornito et al. (2016), ecology Bascompte (2007), and epidemiology Pastor-Satorras and Vespignani (2001), among others Newman (2003). Understanding the evolutionary dynamics of complex networks, as they adapt their connectivity patterns to achieve diverse goals, is crucial to understanding their long-term stability or other features influencing functional roles and performance Fortuna et al. (2011); Villegas et al. (2020). Notably, amidst the multitude of potential network structures, one organization ubiquitously arises in natural systems: the scale-free architecture Barabási and Albert (1999); Newman (2010); Barabási (2016); Caldarelli et al. (2002).
@@
Scale-free networks manifest a distribution of node connectivities $k$ that decays as a power-law $P(k)\propto k^{-\gamma}$ for large values of $k$ Barabási and Albert (1999); Broido and Clauset (2019). In statistical physics, power-law behavior is the hallmark of scale invariance and scaling behavior, implying no significant characteristic value of the analyzed quantity Newman (2005); Sornette (2006); Muñoz (2018). For scale-free networks, there is no typical scale in the degree of connectivity apart from natural cut-offs. However, the network community has been intrigued by the possibility of discerning, through careful statistical analyses Clauset et al. (2009); Broido and Clauset (2019), including finite-size effects Serafino et al. (2021), whether empirical networks genuinely exhibit bona-fide scale invariance or only appear to. From a theoretical standpoint, addressing this question calls for designing a renormalization group (RG) approach Wilson (1979); Amit and Martin-Mayor (2005); Caldarelli et al. (2024); Kardar (2007). In this framework, a fixed point of the RG transformation denotes a state or system whose characteristics remain unchanged under appropriate scale transformations achieved through iterated coarse-graining. RG fixed points are inherently linked to universal scaling laws governing the system, so systems at the same fixed point share identical scaling features and belong to the same universality class Amit and Martin-Mayor (2005); Kardar (2007). Hence, a key implication of RG theory is the categorization of numerous seemingly disparate physical systems and dynamical models into a relatively compact set of universality classes at criticality, either in or out of thermal equilibrium Hohenberg and Halperin (1977); Ódor (2004). Elucidating the relevant ingredients that give rise to a particular universality class offers valuable insights into the fundamental mechanisms underpinning complex systems and their key features.
@@
Given the lack of a natural Euclidean embedding for complex networks, traditional length-scale transformations using translational invariance were deemed unfeasible until recently Radicchi et al. (2008); Rozenfeld et al. (2007). In particular, small-world effects, characterized by short path lengths between nodes, further complicate block identification or affine transformations Cohen and Havlin (2003); Watts and Strogatz (1998); Rozenfeld et al. (2007, 2010). Nevertheless, various techniques have been proposed to tile the network, exploiting specific scale-invariant properties as a function of the number of links along any shortest path between two nodes Radicchi et al. (2008). These include box-covering techniques Song et al. (2005); Rozenfeld et al. (2010), spectral partitioning Gfeller and De Los Rios (2007), and hyperbolic geometry embeddings, which offer a promising novel approach for understanding particular complex network structures and dynamics Serrano et al. (2008); García-Pérez et al. (2018). However, fully characterizing scale-invariant properties in real networks remains an open challenge.
@@
A novel approach, the Laplacian Renormalization Group (LRG), generalizes statistical mechanics RG to graphs, providing a comprehensive framework for coarse-graining heterogeneous systems in both real and momentum space Villegas et al. (2023a). This approach uses diffusive dynamics on networks to gradually remove the smallest fine-grained structural scales, eliminating the contribution from large eigenvalues. Specifically, the LRG allows for the redefinition of effective time-dependent Laplacian and adjacency matrices at coarser scales. Due to its general validity, it comes natural to explore structural scale invariance within the LRG framework, as evidenced by the observation of a time-independent rate of entropy loss Villegas et al. (2023a); Klemm (2023). This leads to the fundamental question of whether non-Euclidean architectures, like general networks, can be intrinsically scale-invariant. Two key questions arise. Can we identify benchmark classes of network architectures exhibiting genuine scale-invariant behavior and, thus, define an analog to universality classes? Can we detect and quantify this kind of self-similarity in real-world networks? Here, we provide a clear definition and compilation of scale-invariant networks, emphasizing that scale-freeness, i.e., a power law degree distribution, does not imply structural self-similarity, and vice versa. Finally, we analyze actual brain networks and highlight their scale-invariant features.
@@
The introduction of a multi-scale statistical mechanical approach, the LRG, allows the detection of different scales within a network Villegas et al. (2023a, b). The LRG is based on the time-evolution operator $e^{-\tau\hat{L}}$ of the diffusion or heat equation, where $\hat{L}=\hat{D}-\hat{A}$ is the Laplacian operator, $\hat{A}$ is the adjacency matrix, and $\hat{D}$ is the degree matrix Villegas et al. (2022). Using this operator and denoting the Laplacian eigenvalues as $\lambda_{i}$ with $i=1,...,N$ (all real and positive Chung (1997); Donetti et al. (2006)), one can define the Laplacian density matrix De Domenico and Biamonte (2016),
@@
and develop a "canonical" description of heterogeneous networks in analogy with statistical mechanics Villegas et al. (2023a, 2022). Here, $\hat{L}$ formally acts as a Hermitian Hamiltonian, $Z(\tau)$ is the partition function, and $\tau$ serves as a control scale parameter akin to the inverse temperature. In this way, as the resolution scale $\tau$ is increased, the contribution of large eigenvalues to $\hat{\rho(\tau)}$revealing fine structure- is progressively removed, allowing for an effective network coarse-graining  [^1] This idea is also at the basis of different existing community detection algorithms Villegas et al. (2023b); Arenas et al. (2006); Donetti and Muñoz (2004); Donetti et al. (2005). Within this formalism, one can compute the network entropy as $S(\tau)=-\mbox{Tr}[\hat{\rho}(\tau)\log\hat{\rho}(\tau)]$, so that $S(\tau)=\tau\langle\lambda\rangle_{\tau}+\ln Z(\tau)$, where $\langle\lambda\rangle_{\tau}\equiv\langle\hat{L}\rangle_{\tau}=\nicefrac{{\stackrel{{\scriptstyle[}}{{i}}=1]{N}{\sum}\lambda_{i}e^{-\tau\lambda_{i}}}}{{Z(\tau)}}$. In particular, $S$ runs from $S=\ln N$ at $\tau=0$, the segregated (high temperature) regime to $S=0$ at $\tau\rightarrow\infty$, the integrated (low temperature) regime. More specifically, one can define the network entropic susceptibility or "heat capacity" Villegas et al. (2022)
$C(\tau)\equiv-\frac{dS}{d\log\tau},$ (2)
@@
which describes the rate of entropy loss —i.e., the rate at which the complexity of the network shrinks upon coarse-graining— or the rate of information acquired about the network structure during diffusion dynamics at scale $\tau$. Recalling statistical physics, peaks of $C$ (diverging in the infinite size limit) are associated with structural phase transitions. Thus, we can analyze $C$ at varying $\tau$ to investigate the network multi-scale organization, detecting scales where entropy changes more significantly due to structural transitions. This description permitted, e.g., the detection of the network information core and its related structural/diffusive transitions Villegas et al. (2022) and led to naturally extending the RG to heterogeneous networks Villegas et al. (2023a). Here, it allows us to define informationally scale-invariant networks as graphs whose entropy-loss rate $C(\tau)$ takes a constant value $C_{0}>0$ across scales or, at least, within a sufficiently broad diffusion-time interval, thus being (exactly or approximately) scale invariant.
@@
We demonstrate that this definition of informational scale-invariance holds if and only if the Laplacian spectral density follows $w(\lambda)\sim\lambda^{\gamma}$, considering the spectral density as a continuum distribution in the infinite network-size limit. Using Eq.(2) and the definition of $S(\tau)$ one derives $C(\tau)=-\tau^{2}\frac{d\langle\lambda\rangle_{\tau}}{d\tau}$, and setting $C(\tau)=C_{0}$ as a constant, one gets $\langle\lambda\rangle_{\tau}=C_{0}/\tau$. The only solution to this equation is a Laplacian spectral density $w(\lambda)\sim\lambda^{\gamma}$, which allows expressing $C_{0}$ as a function of the exponent $\gamma$, specifically $C_{0}=\gamma+1$. This implies that macroscopic properties —stemming from vanishing Laplacian eigenvalues— of a network with a power-law spectrum remain invariant under LRG scale transformations. Thus, network scale-invariance shifts from a power-law degree distribution to a power-law in the spectral density. Note that the exponent of the Laplacian spectral density, $\gamma$, can be related to the spectral dimension, $d_{s}$, which is a global property of the graph, related, e.g., to the infrared singularity of the Gaussian process Burioni and Cassi (1996) and that has been shown to provide a robust generalization of the standard dimension for networks Cassi (1992); Burioni and Cassi (1996). As $d_{s}/2=\gamma+1$ Burioni and Cassi (1996), one concludes that $C_{0}$ is constant and equals half of the graph’s spectral dimension: $C_{0}=d_{s}/2$. Thus, measuring the plateau value of the heat capacity of a scale-invariant network effectively determines its dimensionality.
@@
For finite-sized networks, scale invariance may only be approximate and confined to a finite-scale interval. This means that $C(\tau)$ cannot remain constant across all scales: it decays from its plateau value $C_{0}$ to 00 for sufficiently large values of $\tau$. To estimate this cut-off, we have to consider that the smallest non-zero eigenvalue $\lambda_{2}$ of $\hat{L}$, also called "spectral gap" or "Fiedler eigenvalue" $\lambda_{F}$, Fiedler (1989); Chung et al. (2003), is the one providing the slowest decaying contribution in time to Eq.(1), therefore determining the asymptotic time decay of $C(\tau)$. Starting from Eq.(2) and imposing that the Laplacian spectrum $\omega(\lambda)$ integrates to $N$, we get $\lambda_{F}\sim N^{\nicefrac{{-2}}{{d_{s}}}}$ for large $N$ Das (2004); Dorogovtsev and Mendes (2022). From this, it follows (see Supplemental Material, SM SM ),
${\small C\left(\frac{\alpha}{\lambda_{F}}\right)\simeq\frac{\Gamma\left(\frac{d_{s}}{2}\right)\Gamma\left(\frac{d_{s}}{2}+2,\alpha\right)-\Gamma^{2}\left(\frac{d_{s}}{2}+1,\alpha\right)}{\Gamma^{2}\left(\frac{d_{s}}{2}\right)},}$ (3)
@@
The LRG method relies on a “heat-like" diffusion process defined by $\hat{L}$. However, we wonder whether the results obtained for scale-invariance and graph dimensionality remain valid when using the Laplacian random-walk operator, $\hat{L}_{RW}=\hat{D}^{-1}\hat{L}$ instead of $\hat{L}$. $\hat{L}_{RW}$ describes the time-discrete dynamics of a RW moving from a node to a neighbor with uniform probability and represents the transition matrix for RW dynamics on a graph Burioni and Cassi (2005); Chung (1997). Although $\hat{L}_{RW}$ is not symmetric, it is equivalent to the symmetric operator $\hat{L}_{sym}=D^{1/2}\hat{L}_{RW}\hat{D}^{-1/2}$ Chung (1997) with all its eigenvalues real and satisfying $0\leq\mu_{i}\leq 2$. We can, thus, reformulate the LRG for the RW dynamics by substituting $\hat{L}$ with $\hat{L}_{sym}$, renaming the related heat capacities as $C_{L}$ and $C_{RW}$, respectively. In fact, $\hat{L}_{RW}$ provided the original definition of the RW spectral dimension, $d_{s}^{RW}$, related to the first-return time distribution of the random-walker $P_{0}(t)$ Burioni and Cassi (1996); Cassi and Fabbian (1999) by the scaling $P_{0}(t)\propto t^{-d_{s}^{RW}/2}$. This extends the behavior on $d-$dimensional lattices where $P_{0}(t)\propto t^{-d/2}$ Alexander and Orbach (1982). While $\hat{L}$ or $\hat{L}_{RW}$ strictly coincide in regular networks like lattices, both operators unravel heterogeneous network structures in different ways Villegas et al. (2023b) and, to our knowledge, there is no proof that $d_{s}=d_{s}^{RW}$ for generic heterogeneous networks.
LRG non-trivial fixed points.
@@
Subsequently, we tackle the following lingering questions: What are the current non-trivial fixed points of the LRG, i.e., families or "universality classes" of scale-invariant networks? Are scale-invariant networks necessarily scale-free? Do results depend on the choice of the operator?
@@
Let us remark that regular lattices, with a degree distribution $P(\kappa)=\delta(\kappa-\kappa_{0})$, represent the simplest case of scale-invariant structures lacking scale-free properties (see SM SM ). However, our main focus is categorizing heterogeneous and stochastic scale-invariant networks.
@@
Figure 1: Trees. Heat capacity versus diffusion time for: (a)-(b) Random trees and (c)-(d) Barabasi-Albert networks with $m=1$. We have used $\hat{L}$ in (a) and (c) and $\hat{L}_{RW}$ in (b) and (d). Insets show the scaling of $\tau_{F}$ as a function of the network size $N$. Orange dashed lines and solid black lines represent the theoretical expectation, $d_{s}^{RT}=\nicefrac{{4}}{{3}}$ and $d_{s}^{BA}=2$. All curves have been averaged over $10^{3}$ independent realizations. Different colors represent different sizes $N=\{1,2.5,4,8,16\}\times 10^{3}$.
@@
The first category of self-similar networks comprises trees, i.e., connected loopless networks. It is known that the spectral dimension of random trees, with minimal branching ratio $b_{min}=1$, depends upon the first two moments of the degree distribution $P(\kappa)$ Donetti and Destri (2004); Burda et al. (2001). Specifically, $d_{s}=4/3$ if $\langle\kappa^{2}\rangle$ is finite, while the problem remains open when $\langle\kappa^{2}\rangle$ diverges Donetti and Destri (2004); Destri and Donetti (2002); Dorogovtsev et al. (2003). Here, we examine specific cases within these two classes, with the lenses of LRG: ordinary random trees (RT) and Barabási-Albert (BA) networks where new nodes attach preferentially to existing ones, forming $m=1$ edges Barabási and Albert (1999). As reported in Fig.1(a), $C_{L}$ shows a plateau corresponding to the theoretically known spectral dimension for random trees, $d_{s}^{RT}=\nicefrac{{4}}{{3}}$ Destri and Donetti (2002) and follows the finite-size scaling condition, Eq.(3). Also, we observe that $\hat{L}_{RW}$ does not alter this value (see Fig.1(b)), despite specific differences in the shape of $C$, as the local peak at short times in Fig.1(b).
@@
Instead, BA networks with $m=1$ are trees with diverging $\langle\kappa^{2}\rangle$. As shown in Fig.1(c), they have a constant heat capacity $C_{L}$ corresponding to a spectral dimension $d_{s}=2$. Although $C_{RW}$ grows in the intermediate regime, it also becomes flat for asymptotic times and large network sizes (see Fig.1), confirming scale-invariance with $d_{s}^{RW}=d_{s}=2$. This same spectral dimension is also obtained for Bethe lattices with coordination number $z\geq 3$ (see Erzan and Tuncer (2020) and SM SM ). Furthermore, BA networks with $m>1$ show no sign of scale invariance (see SM SM ), proving that the archetypes of scale-free networks are not scale-invariant. This prevents preferential attachment from generating self-similar networks with $d_{s}>2$.
@@
Second, we consider networks that, contrary to trees, have a non-vanishing clustering coefficient. In particular, the Dorogovstev-Golstev-Mendes Dorogovtsev et al. (2002) networks and their generalization: (u,v)-flowers Dorogovtsev et al. (2002); Rozenfeld et al. (2007, 2009). These deterministic fractal-like structures grow by iteratively replacing the link between two nodes with two paths of lengths $u$ and $v$ (see SM SM ). Diameter-based analyses, such as calculating the Hausdorff dimension of these networks lead to an apparent paradox: all $(1,v)$ flowers are infinite-dimensional and exhibit anomalous scaling functions Rozenfeld et al. (2007); Dorogovtsev et al. (2002); Peng and Agliari (2017). This paradox stems from the lack of network embedding in an Euclidean space. Conversely, as shown in Figs.2(a), (b) and SM SM , all flowers exhibit well-defined spectral dimensions Rozenfeld et al. (2007), $d_{s}=\log(\emph{u}+\emph{v})/\log(\emph{u}\emph{v})$ changing from $d_{s}\approx 3.17$ to $d_{s}=1$.

### v3

@@ Abstract
Scale invariance profoundly influences the dynamics and structure of complex systems, spanning from critical phenomena to network architecture. Here, we propose a precise definition of scale-invariant networks by leveraging the concept of a constant entropy-loss rate across scales in a renormalization-group coarse-graining setting. This framework enables us to differentiate between scale-free and scale-invariant networks, revealing distinct characteristics within each class. Furthermore, we offer a comprehensive inventory of genuinely scale-invariant networks, both natural and artificially constructed, demonstrating, e.g., that the human connectome exhibits notable features of scale invariance. Our findings open new avenues for exploring the scale-invariant structural properties crucial in biological and socio-technological systems.
@@
The network paradigm captures essential attributes of real-world complex systems, offering a natural framework for studying entangled interconnected systems across disciplines like neuroscience Fornito et al. (2016), ecology Bascompte (2007), and epidemiology Pastor-Satorras and Vespignani (2001), among others Newman (2003). Understanding the evolutionary dynamics of complex networks, as they adapt their connectivity patterns to achieve diverse goals, is crucial to understanding their long-term stability or other features influencing functional roles and performance Fortuna et al. (2011); Villegas et al. (2020). Notably, amidst the multitude of potential network structures, one organization ubiquitously arises in natural systems: the scale-free architecture Barabási and Albert (1999); Newman (2010); Barabási (2016); Caldarelli et al. (2002).
@@
Scale-free networks manifest a distribution of node connectivities $k$ that decays as a power-law $P(k)\propto k^{-\gamma}$ for large values of $k$ Barabási and Albert (1999); Broido and Clauset (2019). In statistical physics, power-law behavior is the hallmark of scale invariance and scaling behavior, implying no significant characteristic value of the analyzed quantity Newman (2005); Sornette (2006); Muñoz (2018). For scale-free networks, there is no typical scale in the degree of connectivity apart from natural cut-offs. However, the network community has been intrigued by the possibility of discerning, through careful statistical analyses Clauset et al. (2009); Broido and Clauset (2019), including finite-size effects Serafino et al. (2021), whether empirical networks genuinely exhibit bona-fide scale invariance or only appear to. From a theoretical standpoint, addressing this question calls for designing a renormalization group (RG) approach Wilson (1979); Amit and Martin-Mayor (2005); Caldarelli et al. (2024); Kardar (2007). In this framework, a fixed point of the RG transformation denotes a state or system whose characteristics remain unchanged under appropriate scale transformations achieved through iterated coarse-graining and rescaling. RG fixed points are inherently linked to universal scaling laws governing the system, so systems at the same fixed point share identical scaling features and belong to the same universality class Amit and Martin-Mayor (2005); Kardar (2007). Hence, a key implication of RG theory is the classification of numerous seemingly disparate physical systems and dynamical models into a relatively compact set of universality classes at criticality, whether in or out of thermal equilibrium Hohenberg and Halperin (1977); Ódor (2004). Elucidating the relevant ingredients that give rise to a particular universality class offers valuable insights into the fundamental mechanisms underpinning complex systems and their key features.
@@
Given the lack of a natural Euclidean embedding for complex networks, traditional length-scale transformations using translational invariance were deemed unfeasible until recently Radicchi et al. (2008); Rozenfeld et al. (2007). In particular, small-world effects, characterized by short path lengths between nodes, further complicate block identification or affine transformations in networks Cohen and Havlin (2003); Watts and Strogatz (1998); Rozenfeld et al. (2007, 2010). Nevertheless, various techniques have been proposed to tile networks Radicchi et al. (2008). These include box-covering techniques Song et al. (2005); Rozenfeld et al. (2010), spectral partitioning Gfeller and De Los Rios (2007), and hyperbolic geometry embeddings, which offer a promising novel approach for understanding particular complex network structures and dynamics Serrano et al. (2008); García-Pérez et al. (2018). However, fully characterizing scale-invariant properties in real networks remains an open challenge.
@@
A novel approach, the Laplacian Renormalization Group (LRG), generalizes RG to graphs, providing a comprehensive framework for coarse-graining heterogeneous systems in real and momentum space Villegas et al. (2023a). This approach uses diffusive dynamics on networks to gradually remove the smallest fine-grained structural scales, by eliminating the contribution from large Laplacian eigenvalues. Specifically, the LRG allows for the redefinition of effective Laplacian and adjacency matrices at coarse-grained scales. Due to its general validity, it comes natural to explore structural scale invariance within the LRG framework, as evidenced by the observation of a time-independent rate of entropy loss Villegas et al. (2023a); Klemm (2023). This leads to the fundamental question of whether non-Euclidean architectures, like general networks, can be intrinsically scale-invariant. Two key questions arise. Can we identify benchmark classes of networks exhibiting genuine scale-invariant behavior and, thus, define an analog to universality classes? Can we detect and quantify this kind of self-similarity in real-world networks? Here, we provide a clear definition and compilation of scale-invariant networks, emphasizing that scale-freeness, i.e., a power law degree distribution, does not imply structural self-similarity, and vice versa. Finally, we analyze real brain networks and highlight their scale-invariant features.
@@
Using a statistical mechanical approach, the LRG, allows for the detection of relevant scales in weighted and undirected networks Villegas et al. (2023a, b). The LRG is based on the time-evolution operator at time $\tau$, $e^{-\tau\hat{L}}$ of the diffusion or heat equation, where $\hat{L}=\hat{D}-\hat{A}$ is the Laplacian operator, $\hat{A}$ the adjacency matrix, and $\hat{D}$ the diagonal degree matrix Villegas et al. (2022). Using this and denoting the Laplacian eigenvalues as $\lambda_{i}$ with $i=1,...,N$ (all real and positive Chung (1997); Donetti et al. (2006)), one can define the Laplacian density matrix De Domenico and Biamonte (2016),
@@
and develop a "canonical" description of heterogeneous networks, fully analogous to that in statistical mechanics Villegas et al. (2023a, 2022). Note that, $\hat{L}$ formally acts as a Hermitian Hamiltonian, $Z(\tau)$ is the partition function, and $\tau$ serves as a control scale parameter analogous to the inverse temperature. In this way, as the resolution scale $\tau$ is increased, the contribution of large eigenvalues to $\hat{\rho(\tau)}$ —revealing fine structure— is progressively removed, allowing for an effective network coarse-graining  [^1] This idea is also at the basis of different existing community detection algorithms Villegas et al. (2023b); Arenas et al. (2006); Donetti and Muñoz (2004); Donetti et al. (2005). Thus, one can compute the network entropy as $S(\tau)=-\mbox{Tr}[\hat{\rho}(\tau)\log\hat{\rho}(\tau)]$, so that $S(\tau)\sim\tau\langle\lambda\rangle_{\tau}+\ln Z(\tau)$, where $\langle\lambda\rangle_{\tau}\equiv\langle\hat{L}\rangle_{\tau}=Tr[\hat{\rho}(\tau)\hat{L}]=\nicefrac{{\stackrel{{\scriptstyle[}}{{i}}=1]{N}{\sum}\lambda_{i}e^{-\tau\lambda_{i}}}}{{Z(\tau)}}$. In particular, $S$ runs from $S=\ln N$ at $\tau=0$, the segregated regime to $S=0$ at $\tau\rightarrow\infty$, the fully integrated regime. Moreover, one can define the network entropic susceptibility or "heat capacity" Villegas et al. (2022) as,
$C(\tau)\equiv-\frac{dS}{d\log\tau}=-\tau^{2}\frac{d\langle\lambda\rangle_{\tau}}{d\tau},$ (2)
@@
describing the rate of entropy loss (at which the complexity of the network shrinks upon coarse-graining) or the rate of information acquired about the network structure during diffusion dynamics at scale $\tau$. In analogy with statistical physics, peaks of $C$ (diverging in the infinite size limit) are associated with phase transitions. Thus, we can analyze $C$ at varying $\tau$ to investigate the network multi-scale organization, detecting scales where entropy changes more significantly due to structural transitions. This description allowed, for example, the detection of the network information core and its associated structural and diffusive transitions Villegas et al. (2022), and led to the natural extension of RG to heterogeneous networks Villegas et al. (2023a). Here, it allows us to define informationally scale-invariant networks as graphs whose entropy-loss rate $C(\tau)$ takes a constant value $C_{0}>0$ across scales or, at least, within a sufficiently broad diffusion-time interval, thus being (exactly or approximately) scale invariant.
@@
We demonstrate that this definition of informational scale-invariance holds if the Laplacian spectral density follows $P(\lambda)\sim\lambda^{\gamma}$, considering it as a continuum distribution in the infinite network-size limit. Using Eq.(2) and setting $C(\tau)=C_{0}$ as a constant, one gets $\langle\lambda\rangle_{\tau}=C_{0}/\tau$. The only solution to this equation is a Laplacian spectral density $P(\lambda)\sim\lambda^{\gamma}$, which allows us to express $C_{0}$ as a function of the exponent $\gamma$, specifically $C_{0}=\gamma+1$. This implies that macroscopic properties —stemming from small Laplacian eigenvalues— of a network with a power-law spectrum remain invariant under LRG transformations. Thus, the concept of network scale-invariance shifts from a power-law degree distribution to a power-law in the spectral density. Note that the exponent of the Laplacian spectral density, $\gamma$, can be related to the spectral dimension, $d_{s}$, which is a global property of the graph, related, e.g., to the infrared singularity of the Gaussian process Burioni and Cassi (1996) and that has been shown to provide a robust generalization of the standard concept of dimension for networks Cassi (1992); Burioni and Cassi (1996). As $d_{s}/2=\gamma+1$ Burioni and Cassi (1996), one concludes that $C_{0}$ is constant and equals half of the graph’s spectral dimension: $C_{0}=d_{s}/2$. Thus, measuring the plateau value of the heat capacity of a scale-invariant network effectively determines its dimensionality.
@@
For finite-sized networks, scale invariance may only be approximate and confined to a finite-scale interval. This means that $C(\tau)$ cannot remain constant across all scales: it decays from its plateau value $C_{0}$ to 00 for sufficiently large values of $\tau$. To estimate this cut-off, we have to consider that the smallest non-zero eigenvalue $\lambda_{2}$ of $\hat{L}$, also called "spectral gap" or "Fiedler eigenvalue" $\lambda_{F}$, Fiedler (1989); Chung et al. (2003), is the one providing the slowest decaying contribution in time to Eq.(1), therefore determining the asymptotic time decay of $C(\tau)$. Starting from Eq.(2) and imposing that the Laplacian spectrum $P(\lambda)$ integrates to $N$, one obtains $\lambda_{F}\sim N^{\nicefrac{{-2}}{{d_{s}}}}$ for large $N$ Das (2004); Dorogovtsev and Mendes (2022). From this, it follows that for large times $\tau_{F}(\alpha)=\alpha/\lambda_{F}$, where $\alpha\gg 1$ is a free parameter (see Supplemental Material, SM SM )
${\small C\left(\tau_{F}(\alpha)\right)\simeq\frac{\Gamma\left(\frac{d_{s}}{2}\right)\Gamma\left(\frac{d_{s}}{2}+2,\alpha\right)-\Gamma^{2}\left(\frac{d_{s}}{2}+1,\alpha\right)}{\Gamma^{2}\left(\frac{d_{s}}{2}\right)},}$ (3)
@@
Laplacian Random-Walk RG. The LRG method relies on a “heat-like" diffusion process defined by $\hat{L}$. However, we wonder whether the results obtained for scale-invariance and graph dimensionality remain valid when using the Laplacian random-walk operator, $\hat{L}_{RW}=\hat{D}^{-1}\hat{L}$ instead of $\hat{L}$. $\hat{L}_{RW}$ describes the time-discrete dynamics of a RW moving from a node to a neighbor with uniform probability and is related to the transition matrix for RW dynamics on a graph Burioni and Cassi (2005); Chung (1997). Although $\hat{L}_{RW}$ is not symmetric, it is similar to the symmetric operator $\hat{L}_{sym}=D^{1/2}\hat{L}_{RW}\hat{D}^{-1/2}$ Chung (1997) with all its eigenvalues real and satisfying $0\leq\mu_{i}\leq 2$. We can, thus, reformulate the LRG for the RW dynamics by substituting $\hat{L}$ with $\hat{L}_{sym}$, renaming the related heat capacities as $C_{L}$ and $C_{RW}$, respectively. In fact, $\hat{L}_{RW}$ provided the original definition of the RW spectral dimension, $d_{s}^{RW}$, related to the first-return time distribution of the random-walker $P_{0}(t)$ Burioni and Cassi (1996); Cassi and Fabbian (1999) by the scaling $P_{0}(t)\propto t^{-d_{s}^{RW}/2}$. This generalizes the behavior on $d-$dimensional lattices where $P_{0}(t)\propto t^{-d/2}$ Alexander and Orbach (1982). While $\hat{L}$ or $\hat{L}_{RW}$ strictly coincide in regular networks like lattices, both operators unravel heterogeneous network structures in different ways Villegas et al. (2023b) and, to our knowledge, there is no proof that $d_{s}=d_{s}^{RW}$ for generic heterogeneous networks.
@@
LRG non-trivial fixed points. We now tackle the following lingering questions: What are the current non-trivial fixed points of the LRG, i.e., families or "universality classes" of scale-invariant networks? Are scale-invariant networks necessarily scale-free? Do results depend on the choice of the operator?
@@
Let us remark that regular lattices, with a degree distribution $P(\kappa)=\delta(\kappa-\kappa_{0})$, are the simplest case of scale-invariant structures lacking scale-free properties (see SM SM ). However, our main focus is the categorization of heterogeneous and stochastic scale-invariant networks.
@@
Figure 1: Trees. Heat capacity as a function of the resolution scale $\tau$: (a)-(b) RTs and (c)-(d) BA networks with $m=1$. We use $\hat{L}$ in (a) and (c) and $\hat{L}_{RW}$ in (b) and (d). Insets show the scaling of $\tau_{F}$ as a function of network size $N$. Orange dashed lines and solid black lines represent the scaling for the theoretical expectation, $d_{s}^{RT}=\nicefrac{{4}}{{3}}$ and $d_{s}^{BA}=2$. All curves are averages over $10^{3}$ independent realizations. Different colors represent different sizes $N=\{1,2.5,4,8,16\}\times 10^{3}$.
@@
The first category of self-similar networks comprises trees, i.e., connected loopless networks. It is known that the spectral dimension of random trees, with minimal branching ratio $b_{min}=1$, depends upon the first two moments of the degree distribution $P(\kappa)$ Donetti and Destri (2004); Burda et al. (2001). Specifically, $d_{s}=4/3$ if $\langle\kappa^{2}\rangle$ is finite, while the problem remains open when $\langle\kappa^{2}\rangle$ diverges Donetti and Destri (2004); Destri and Donetti (2002); Dorogovtsev et al. (2003). Here, we examine specific cases within these two classes, using the lenses of LRG: ordinary random trees (RT) and scale-free Barabási-Albert (BA) networks where new nodes attach preferentially to existing ones, forming $m=1$ edges Barabási and Albert (1999). As reported in Fig.1(a), $C_{L}$ shows a plateau corresponding to the theoretically known spectral dimension for RTs, $d_{s}^{RT}=\nicefrac{{4}}{{3}}$ Destri and Donetti (2002) and obeys the finite-size scaling condition, Eq.(3) for various network sizes. We also observe that $\hat{L}_{RW}$ does not alter this value (see Fig.1(b)), despite specific differences in the $C$ shape, as the local peak at short times in Fig.1(b).
@@
Instead, scale-free BA networks with $m=1$ are trees with diverging $\langle\kappa^{2}\rangle$. As shown in Fig.1(c), they reach a constant heat capacity $C_{L}$ corresponding to a spectral dimension $d_{s}=2$. $C_{RW}$ grows in the intermediate regime but becomes flat for asymptotic times and large network sizes (see Fig.1), confirming scale-invariance with $d_{s}^{RW}=d_{s}=2$. This same spectral dimension is also obtained for Bethe lattices with coordination number $z\geq 3$ (see Erzan and Tuncer (2020) and SM SM ). Furthermore, BA networks with $m>1$ show no sign of scale invariance (see SM SM ), proving that the archetypes of scale-free networks are not scale-invariant. This prevents preferential attachment from generating self-similar networks with $d_{s}>2$.
@@
Second, we consider networks that, contrary to trees, have a non-vanishing clustering coefficient. In particular, Dorogovtsev-Goltsev-Mendes Dorogovtsev et al. (2002) networks and their generalization: (u,v)-flowers Dorogovtsev et al. (2002); Rozenfeld et al. (2007, 2009). These deterministic fractal-like structures grow by iteratively replacing the link between two nodes with two paths of lengths $u$ and $v$, respectively (see SM SM ). Diameter-based analyses, such as calculating the Hausdorff dimension of these networks lead to an apparent paradox: all $(1,v)$ flowers are infinite-dimensional and exhibit anomalous scaling functions Rozenfeld et al. (2007); Dorogovtsev et al. (2002); Peng and Agliari (2017). This paradox stems from the lack of network embedding in an Euclidean space. Conversely, as shown in Figs.2(a/b) and in the SM SM , all flowers exhibit well-defined spectral dimensions Rozenfeld et al. (2007), $d_{s}=2\log(\emph{u}+\emph{v})/\log(\emph{u}\emph{v})$ changing from $d_{s}\approx 3.17$ to $d_{s}=1$.